Find three consecutive odd integers such that the sum of the first and third integers is twice the second integer.
Answer by nvela(1)

n + (n + 4) = 2(n + 2)
2n + 4 = 2n + 4
2n (-2n) + 4 = 2n (-2n) + 4
4=4
1
n + 2 = 1 + 2 = 3
n + 4 = 1 + 4 = 5
Answer: 1,3,5